# Low-hallucination Synthetic Captions for Large-Scale Vision-Language Model Pre-training

### v1

@@ 5 Conclusion and Limitation
5.1 Conclusion
In this work, we tackle the issue of synthesizing captions that are rich in knowledge and exhibit low levels of hallucination. We present a novel pipeline for high-quality synthetic captions, featuring two effective techniques: continuous DPO and knowledge-enriching SFT. We elucidate the beneficial role of DPO in alleviating hallucinations within image captions, along with its scaling law within a certain scope. Furthermore, we propose an approach to continually improve model performance once the scaling law ceases to be effective. Experimental results demonstrate that Hunyuan-Recap100M outperforms original alt-text and other synthetic captions in the pre-training of vision language models by a large margin. We will make Hunyuan-Recap100M publicly available to promote the development of multimodal related fields in the open source community.
Limitations.
Like most of the existing work on vision language models, the entire project consumed a lot with high electricity costs, although we only tested a $7B$ size model on $100M$ training data. Due to limited computing power, we did not scale up the use of recaption data to a large extent in this paper. We leave this work for future research. We also did not verify the effectiveness of the large-scale recaption data we generated for training text-to-image models from scratch. This will also be considered as future work.
Potential Risks.
The original image-text pairs are mainly derived from open source datasets. Although we have taken measures to detoxify to a great extent, there are still potential risks, especially in the field of image generation, such as fake portraits on social media (Hill and White, 2020), which is a common potential risk in image generation research.
@@ Appendix A Loss of Vision-Language Model Training
To address the issue of incomparability of loss arising from diverse data distributions, we examine the convergence of pre-trained models that are trained on different datasets but are evaluated on the same supervised fine-tuning data, as illustrated in Figure 5. In particular, it can be observed that models trained with Hunyuan-Recap100M display a better convergence performance when dealing with the same data from downstream tasks. Specifically, these models exhibit lower loss values at both the starting point, with a loss value of $0.925$, and the convergence point, where the loss value is $0.48$.
Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B or Hunyuan-Recap100M. Note that all pre-trained models use the same amount of data (20M), and at the same time, the three curves are evaluated on exactly the same supervised fine-tuning data.
@@ Appendix B Examples of Text-to-Image Generations
In this section, we present a series of examples of text-to-image generation in Figure 6. The synthetic captions corresponding to the text concepts in Figure 6 can be found in Table 6. From these examples, we can see that the HunyuanDiT model does not respond to a few concepts, such as abacus and Yueqin, and its depictions of some other concepts are also inaccurate, like jellyfish and acerola. After tuning with Hunyuan-Recap100M, the model’s understanding of these concepts becomes more precise, and its generated image style is also more realistic. The situations demonstrated in these cases are consistent with the reduction in FID shown in Table 5, highlighting the wide coverage of concepts perceived by our data. This is corroborated by the significant improvement in the performance of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks, as presented in Table 4.
Text Image HunyuanDiT HunyuanDiT wRecap HunyuanDiT wFT
@@
Figure 6: Demonstration of text-to-image cases. Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "HunyuanDiT", the generation effect of the original open-source model with our Synthetic texts as "HunyuanDiT wRecap", and the effect of the model fine-tuned with our recaption data as "HunyuanDiT wFT".

### v2

@@ 5 Conclusion and Limitation
This work addresses the challenge of synthesizing low-hallucination, knowledge-rich image captions. We introduce a novel pipeline for generating such high-quality synthetic captions employing two key techniques: continuous DPO (CDPO) and knowledge-enriching SFT. We demonstrate that DPO effectively mitigates hallucinations in captions and analyze its scaling behavior at certain data scales. Furthermore, we present the CDPO approach designed to sustain performance improvements beyond the point where DPO’s scaling plateaus. Experimental results show that Hunyuan-Recap100M data significantly outperform original alt-text and other synthetic caption datasets for pre-training vision language models. To foster the development of multimodal research within the open source community, we will make Hunyuan-Recap100M publicly available. See Section 6 and Section 7 for details on limitations and risks.
6 Limitations
Training large vision-language models is inherently computationally intensive and resource-demanding. Specifically, pre-training the 7B-sized VLM model using the complete Hunyuan-Recap100M dataset required substantial $15,360$ H20 GPU compute hours. Due to limited computational resources, this work does not explore the full potential of scaling up the use of recaption data. Furthermore, we did not evaluate the effectiveness of using our generated large-scale recaption data to train text-to-image models from scratch. Addressing these limitations by conducting experiments with larger data scales and exploring T2I training from scratch are important avenues for future research.
7 Potential Risks
The original image-text pairs are primarily derived from open-source datasets. While we have implemented substantial measures to filter out undesirable content, potential risks remain. These risks are particularly salient in the field of image generation, exemplified by issues such as the creation of fake portraits for social media (Hill and White, 2020), which is a recognized challenge in this research area.
@@ Appendix A Loss of Vision-Language Model Training
To address the issue of loss value incomparability arising from diverse pre-training data distributions, we examine the convergence behavior of pre-trained models when fine-tuned on a common supervised dataset [Wang et al., 2023, Li et al., 2024a], as illustrated in Figure 5. Figure 5 shows that models trained with Hunyuan-Recap100M data exhibit superior convergence. Quantitatively, these models converge to a lower final loss of $0.4796$ compared to models pre-trained on other datasets.
Figure 5: Loss curve of vision-language model pre-trained with alt-text, Recap-DataComp-1B, Qwen2-VL-Zeroshot or Hunyuan-Recap100M. To ensure a fair comparison, all pre-trained models analyzed here utilized the same amount of data ($20M$) for pre-training. Furthermore, the four convergence curves presented are obtained by fine-tuning these models on an identical supervised dataset.
@@ Appendix B Examples of Text-to-Image Generations
This section presents examples of text-to-image generation in Figure 6; the corresponding text prompts are provided in Table 6. Observation of these examples reveals that the original Hunyuan-DiT model struggles to accurately render certain concepts (e.g. abacus, Yueqin) and produces less accurate depictions for others (e.g. jellyfish, acerola). However, after fine-tuning with the Hunyuan-Recap100M data, the model’s understanding of these concepts becomes more precise, resulting in generated images with a more realistic style. These qualitative improvements are consistent with the quantitative results shown in Table 5, such as the observed reduction in FID. This consistency highlights how the comprehensive concept coverage in our data enhances the model’s perceptual fidelity, which is further corroborated by the significant performance improvement of the VLM trained with Hunyuan-Recap100M on numerous object perception tasks (Table 4).
Text Image Hunyuan-DiT Hunyuan-DiT wRecap Hunyuan-DiT wFT
@@
Figure 6: Demonstration of text-to-image cases. Here, the original instruction for image generation is denoted as "Text", the original picture as "Image", the generation effect of the original open-source model as "Hunyuan-DiT", the generation effect of the original open-source model with our Synthetic texts as "Hunyuan-DiT wRecap", and the effect of the model fine-tuned with our recaption data as "Hunyuan-DiT wFT".